OpenStudy (anonymous):

How do the graphs of the functions f(x)=x^2-5 and g(x)=x^2+4 relate to each other? A. The graph of f(x) is is 9 units to the left of the graph of g(x). B. The graph of f(x) is 1 unit below the graph of g(x). C. The graph of f(x) is 9 units below the graph of g(x). D. The graph of f(x) is 1 unit to the left of the graph of g(x). 1st correct answerer gets fan + medal!

OpenStudy (anonymous):

I' thinking it is C, but I graphed it and it doesn't look right. Any help?

jimthompson5910 (jim_thompson5910):

The -5 on f(x) = x^2 - 5 shifts x^2 five units down The +4 on g(x)=x^2+4 shifts x^2 four units up

jimthompson5910 (jim_thompson5910):

So their distance is 4 - (-5) = 4 + 5 = 9 units

jimthompson5910 (jim_thompson5910):

Making f(x) 9 units below g(x) or g(x) 9 units above f(x)

OpenStudy (anonymous):

So it's C right?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks :D